A metal fabricator needs a piece of aluminum in the shape of a right triangle. The length of the hypotenuse if 28 inches. If the length of the longer leg is 5 inches more than the shorter leg what is the length of both legs?
Please help. Thanks.

Answer by Boreal(15235)

shorter leg=x
longer leg=x+5
hypotenuse is 28
x^2+(x+5)^2=28^2=784
x^2+x^2+10x+25=784
2x^2+10x-759=0
x+(1/4)(-10+/- sqrt (100+6072); sqrt (6172)=78.56
x=(1/4)(-10+78.56), only positive root, so x=(1/4)(68.56)=17.14
legs are 17.14 in and 22.14in ANSWER
square everything and 293.78+490.18=783.96 which is within rounding error.